# CALRec: Contrastive Alignment of Generative LLMs For Sequential Recommendation

### v1

@@ Abstract.
Traditional recommender systems such as matrix factorization methods rely on learning a shared dense embedding space to represent both items and user preferences. Sequence models such as RNN, GRUs, and, recently, Transformers have also excelled in the task of sequential recommendation. This task requires understanding the sequential structure present in users’ historical interactions to predict the next item they may like. Building upon the success of Large Language Models (LLMs) in a variety of tasks, researchers have recently explored using LLMs that are pretrained on vast corpora of text for sequential recommendation. To use LLMs in sequential recommendations, both the history of user interactions and the model’s prediction of the next item are expressed in text form. We propose CALRec, a two-stage LLM finetuning framework that finetunes a pretrained LLM in a two-tower fashion using a mixture of two contrastive losses and a language modeling loss: the LLM is first finetuned on a data mixture from multiple domains followed by another round of target domain finetuning. Our model significantly outperforms many state-of-the-art baselines ($+37\%$ in Recall@1 and $+24\%$ in NDCG@10) and systematic ablation studies reveal that (i) both stages of finetuning are crucial, and, when combined, we achieve improved performance, and (ii) contrastive alignment is effective among the target domains explored in our experiments.
@@
copyright: noneccs: Information systems Recommender systemsccs: Computing methodologies Neural networks
@@ 1. Introduction and Motivation
Recommender systems aim to understand users and items and use the learned knowledge to recommend relevant items for future user interactions. Traditional collaborative filtering methods (Goldberg et al., 1992; Sarwar et al., 2001; Koren et al., 2009) map item and user IDs into a common latent embedding space learned from individual user-item interactions, but these embeddings are often static and fail to capture the dynamic nature of user interests. Recent exploration in sequential recommender employs RNN (Hidasi et al., 2016), CNN (Yan et al., 2019) and the transformer (Kang and McAuley, 2018; Sun et al., 2019) to directly learn the time-evolution of user behaviors. The transformer architecture  (Vaswani et al., 2017; Devlin et al., 2019) stands out for its efficient capability of modeling sequences through the self-attention mechanism, which learns the correlation or causal relationships between different tokens and their influence on future tokens. In the context of recommender systems (RecSys), each input position (token) corresponds to a single user activity, typically a user’s interaction (e.g., purchase) with an item  (Kang and McAuley, 2018; Sun et al., 2019; Zhang et al., 2019; Zhou et al., 2020; Hou et al., 2022). Some approaches rely on item IDs, each represented by a unique token, and model a user’s interaction history as a sequence of item ID tokens (Kang and McAuley, 2018; Sun et al., 2019); other efforts combine more sophisticated features, such as the text embedding of item descriptions, to capture richer content information for each item  (Zhang et al., 2019; Zhou et al., 2020).
@@
Motivated by the global paradigm shift towards autoregressive Language Large Models (LLMs) (Achiam et al., 2023; Anil et al., 2023; Tay et al., 2023; Team et al., 2023) with successful applications in various tasks and domains (Dosovitskiy et al., 2021; Arnab et al., 2021; Ramesh et al., 2022; Yu et al., 2021), we propose using generative LLMs as an approach for sequential recommender systems. Pretrained LLMs learn knowledge from large text corpora and are expected to perform reasonably well in general tasks, and performance improves on specific tasks after finetuning with target domain data. Several studies have explored utilizing LLMs in recommender systems (Wu et al., 2023), such as framing sequential recommendation as sentence-completion (Li et al., 2023c), question-answering (Zhang et al., 2023) or ranking problems (Kang et al., 2023).
@@
Despite the impressive language understanding capabilities of pretrained LLMs, they require fine-tuning for highly specific tasks like sequential recommendation. An LLM should be ideally fully fine-tuned (Kang et al., 2023; Whitehouse et al., 2023) to learn to comprehend domain-specific flattened text sequences and the underlying user interest evolution. Similar to most sequence modeling approaches like SASRec (Kang and McAuley, 2018), BERT4REC (Sun et al., 2019), Recformer (Li et al., 2023b), and GPT4Rec (Li et al., 2023c), we employ next-item prediction as the sequential recommendation task. Based on sentence-completion (Li et al., 2023c) or question-answering  (Zhang et al., 2023) styled prompting, our template design is further inspired by the idea of few-shot learning (Brown et al., 2020). Specifically, we prefix user sequence and each item in the user sequence with distinctive text indicators that 1) allow the model to better understand data pattern and text format and 2) increase the LLM response coherence of the input sentence.
We propose the Contrastive Aligned Generative LLM Recommendation (CALRec) framework to adapt an LLM for sequential recommendation, inspired by the effectiveness of contrastive learning (CL) in mapping-based RecSys (Hou et al., 2022; Li et al., 2023b) and CL research in NLP and other areas  (Gao et al., 2021; Li et al., 2023a). CALRec features the following:
@@
Similar to previous work (Kang and McAuley, 2018; Sun et al., 2019; Li et al., 2023b), we leverage the Amazon Review dataset (Ni et al., 2019) and compare our results to open-source state-of-the-art (SotA) baselines. We observed a non-negligible percentage of consecutive identical activities by the same users in the raw datasets which might trivialize the next-item prediction task. We thus train and evaluate the model on deduplicated sequences and propose to adopt ’Last Item Repeater’, a training-free, text-statistic baseline for model comparison (§4 and §6). We also revisit the existing evaluation metrics of the previous works, and propose to report both optimistic and pessimistic metric scores for text-based RecSys (§4). Through comprehensive experiments, we discover that our CALRec method demonstrates superior performance and outperforms existing SotA by a considerable margin for most evaluation metrics. A systematic ablation study indicates the effectiveness of the two-stage training paradigm and our contrastive alignment among the various data categories examined in our experiments.
@@ 2. Related Work
Traditional recommender systems like Collaborative Filtering (Goldberg et al., 1992) learn relevance through user-item co-occurrence. Early collaborative filtering approaches like Matrix Factorization (MF) (Koren et al., 2009) are ID-only where both user and item IDs are embedded in a shared embedding space. Later MF improvements incorporated other attributes (Shi et al., 2014) such as time and location. Nevertheless, these approaches did not capture the sequential nature of user behavior which gave rise to a class of Sequential Recommenders. Similar to MF, initial Sequential Recommenders were ID-only like GRU4Rec (Hidasi et al., 2016), SASRec (Kang and McAuley, 2018), and Bert4Rec (Sun et al., 2019), while follow-up work focused on incorporating additional attributes beyond the IDs, with methods such as FDSA (Zhang et al., 2019), S3-Rec (Zhou et al., 2020), and UniSRec (Hou et al., 2022). Due to the challenges (Shi et al., 2014) that IDs pose in real world Recommender Systems, researchers(Hou et al., 2022; Li et al., 2023b) explored removing them in favor of attribute-only models but faced quality challenges because of the valuable memorization characteristics that IDs bring in. Instead, other researchers explored replacing them with IDs derived from attribute embeddings e.g. (Rajput et al., 2023; Singh et al., 2023; Hou et al., 2023) which alleviate some of the infrastructure challenges posed by random IDs in exchange for increased system complexity.
@@
With the rise of pretrained LLM backbones, which are either encoder-only like BERT (Devlin et al., 2019; Li et al., 2023b), encoder-decoder like T5 (Raffel et al., 2020), or decoder-only like PaLM  (Chowdhery et al., 2023; Anil et al., 2023), GPT (Brown et al., 2020), LLama (Touvron et al., 2023), among others, researchers focused on ways to leverage these in the Sequential Recommender tasks primarily to bring along the world knowledge that these models encode in their parameters. Some of these approaches are zero-shot or few-shot (Wang and Lim, 2023; Liu et al., 2023; Hou et al., 2024) where no LLM parameters are updated, others used the LLM to extract item text embeddings also without parameter updates (Hou et al., 2022).
@@
Researchers explored directly finetuning encoder-only, encoder-decoder or decoder-only LLMs to incorporate the Recommender System knowledge into the LLM (Li et al., 2023b; Raffel et al., 2020; Wu et al., 2023; Kang et al., 2023; Zhang et al., 2023; Bao et al., 2023). P5 (Geng et al., 2022) finetuned a pre-trained T5 (Raffel et al., 2020) model to solve 5 common RecSys tasks; Sequential Recommendation is one of them. Unlike P5 which models the Sequential Recommendation task as an ID generation task and its prompt templates does not distinguish between user and item, our work focuses on generating item attribute text and explicitly finetunes the model in a two-tower fashion to explicitly understand the items and their attributes as entries within a user-item interaction sequence. GPT4Rec (Li et al., 2023c) finetunes a pre-trained GPT-2 model to generate the next item titles. Like GPT4Rec, we use BM25 (Robertson et al., 2009) to search for items in the corpus. However, unlike GPT4Rec which directly does supervised finetuning for next item generation for each target domain as a sentence completion task, our work is the first to explore using contrastive loss as training objective for multi-stage finetuning of pre-trained autoregressive language models on sequential recommendation task. This enables the model to generalize to multiple target domains in the final stage of finetuning. We include the list of baseline models we compare against and their details in §4.
@@ 3.1. Preliminaries and Terminology
We assume a set of users $\mathcal{U}$, where a user $u\in\mathcal{U}$ has an interaction sequence that consists of a sequence of $n$ items $(I_{1},I_{2},...,I_{n})$ in chronological order ($n$ can be different for different users). The sequential recommendation task is defined as follows: given the user’s interaction sequence $I_{1:n-1}=(I_{1},I_{2},...,I_{n-1})$, a sequential recommender aims to predict the target item $I_{n}$, the last item in the user’s history. In our study, we use pure text to represent each item and user interaction sequence (§3.2), and we adapt autoregressive LLMs pretrained on large corpora of text for sequential recommendation. Specifically, we pose the task as a sequence-to-sequence generation task where the input is the text description of $I_{1:n-1}$ and the expected output is that of $I_{n}$. The underlying motivation to pose the problem fully in the text domain is to unlock the rich text information in real-world data and directly leverage LLMs’ strong language understanding and reasoning capabilities.
@@ 3.2. Data Format and Template Design
Different from our method, many previous text-based approaches do not specify an item prefix and do not leverage one-off user history suffix (Li et al., 2023b, c; Zhang et al., 2023).
@@ 3.3. Two-Stage Fine-Tuning
Previous work (Hou et al., 2022; Li et al., 2023b) indicates that pretraining a RecSys on a large volume of data (various data categories) provides favorable inductive bias for downstream tasks (a specific and smaller-scale data category for evaluation after further fine-tuning on them). Analogously, based on an LLM already pretrained on large, diverse text corpora, we propose a two-stage fine-tuning framework that consists of Multi-Category Joint Fine-Tuning (Stage I) and Category-Specific Fine-Tuning (Stage II): the latter benefits from the former as a result of transfer learning. Our Stage I fine-tuning allows the model to adapt to the sequential recommendation problem setting and learn data patterns in a category-agnostic fashion. Inspired by Xue et al. (2021), we adopt a sampling rate for each data category proportional to $|\mathcal{U}|^{0.3}$, where $|\mathcal{U}|$ is the size (number of users) of each category, for the purpose of combating the data imbalance issue, preventing a few extra-large categories from dominating the Stage I finetuning set. After deriving the multi-category fine-tuned model, we further adapt the model to each category respectively via Category-Specific Fine-Tuning (Stage II). In both stages, the training objective is the same, introduced in §3.4.
@@ 3.4. Training Objectives
Next Item Generation Objective. The primary objective we use for LLM finetuning is the Next Item (Text) Generation (NIG) Objective. NIG aims to generate the text description of the target item given the text description of the previous items in the user’s history. Denote text-tokenized user sequence as $(t_{1},t_{2},...,t_{l})$, where $l$ is the tokenized sequence length. The first $m$ tokens belong to all the items in the user sequence except for the last item, followed by $l-m$ tokens for the target item. We propose a next item generation objective to adapt an LLM for sequential recommendation:
(1) $\vspace{-5pt}\mathcal{L}_{NIG}=-\mathbb{E}\!\sum_{j=m+1}^{l}\text{log}\,P(t_{j}|t_{1:j-1};\theta)$
where $\theta$ is the whole set of trainable parameters of the chosen LLM.
Auxiliary Contrastive Alignment. One shortcoming of the NIG, is that it deals with the input and output on the token level rather than on the item/user level. Therefore, in addition to the token-level NIG objective, we also investigate auxiliary contrastive objectives that operate on the item/user level. Specifically, we adopt a two-tower training framework where one tower only takes in the target item, and the mean-pooled feature in the final hidden layer is denoted as $\mathbf{v}^{T}$. The second tower takes the whole user sequence and similarly we denote the mean-pooled feature corresponding to the user history sequence as $\mathbf{v}^{U}$ and the feature for the target item conditioned on the user history as $\mathbf{v}^{T|U}$. We experiment with two contrastive losses for user- and item-level alignments following the InfoNCE loss  (Oord et al., 2018), a standard and robust choice in contrastive learning research  (Chen et al., 2020; Gao et al., 2021; Li et al., 2022, 2023a):
@@
Our final training objective is a mixture of $\mathcal{L}_{NIG}$ and contrastive losses as shown in Fig 1:

### v2

@@ Abstract.
Traditional recommender systems such as matrix factorization methods have primarily focused on learning a shared dense embedding space to represent both items and user preferences. Subsequently, sequence models such as RNN, GRUs, and, recently, Transformers have emerged and excelled in the task of sequential recommendation. This task requires understanding the sequential structure present in users’ historical interactions to predict the next item they may like. Building upon the success of Large Language Models (LLMs) in a variety of tasks, researchers have recently explored using LLMs that are pretrained on vast corpora of text for sequential recommendation. To use LLMs for sequential recommendation, both the history of user interactions and the model’s prediction of the next item are expressed in text form. We propose CALRec, a two-stage LLM finetuning framework that finetunes a pretrained LLM in a two-tower fashion using a mixture of two contrastive losses and a language modeling loss: the LLM is first finetuned on a data mixture from multiple domains followed by another round of target domain finetuning. Our model significantly outperforms many state-of-the-art baselines ($+37\%$ in Recall@1 and $+24\%$ in NDCG@10) and our systematic ablation studies reveal that (i) both stages of finetuning are crucial, and, when combined, we achieve improved performance, and (ii) contrastive alignment is effective among the target domains explored in our experiments.
@@
journalyear: 2024copyright: rightsretainedconference: 18th ACM Conference on Recommender Systems; October 14–18, 2024; Bari, Italybooktitle: 18th ACM Conference on Recommender Systems (RecSys ’24), October 14–18, 2024, Bari, Italydoi: 10.1145/3640457.3688121isbn: 979-8-4007-0505-2/24/10ccs: Information systems Recommender systemsccs: Computing methodologies Neural networks
@@ 1. Introduction and Motivation
Recommender systems aim to understand users and items and use the learned knowledge to recommend relevant items for future user interactions. Traditional collaborative filtering methods (Goldberg et al., 1992; Sarwar et al., 2001; Koren et al., 2009) map item and user IDs into a common latent embedding space learned from individual user-item interactions, but these embeddings are often static and fail to capture the dynamic nature of user interests. Recent exploration in sequential recommender employs RNN (Hidasi et al., 2016), CNN (Yan et al., 2019) and the transformer (Kang and McAuley, 2018; Sun et al., 2019) to directly learn the time-evolution of user behaviors. The transformer architecture  (Vaswani et al., 2017; Devlin et al., 2019) stands out for its efficient capability of modeling sequences through the self-attention mechanism, which learns the correlation or causal relationships between different tokens and their influence on future tokens. In a large body of recommender systems (RecSys) research, each input position (token) corresponds to a single user activity, typically a user’s interaction (e.g., purchase) with an item  (Kang and McAuley, 2018; Sun et al., 2019; Zhang et al., 2019; Zhou et al., 2020; Hou et al., 2022). More specifically, some approaches rely on item IDs, each represented by a unique token, and model a user’s interaction history as a sequence of item ID tokens (Kang and McAuley, 2018; Sun et al., 2019); other efforts combine more sophisticated features, such as the text embedding of item descriptions, to capture richer content information for each item  (Zhang et al., 2019; Zhou et al., 2020).
@@
Motivated by the global paradigm shift towards autoregressive Language Large Models (LLMs) (Achiam et al., 2023; Anil et al., 2023; Tay et al., 2023; Team et al., 2023) with successful applications in various tasks and domains (Dosovitskiy et al., 2021; Arnab et al., 2021; Ramesh et al., 2022; Yu et al., 2021), we focus on leveraging generative LLMs for sequential recommender systems. Pretrained LLMs learn knowledge from large text corpora and are expected to perform reasonably well in general tasks, and their performance on specific tasks further improves after fine-tuning with domain-specific data. Several studies have explored utilizing LLMs in recommender systems (Wu et al., 2023), such as framing sequential recommendation as sentence-completion (Li et al., 2023d) or question-answering problems (Zhang et al., 2023).
@@
Despite the impressive language understanding capabilities of pretrained LLMs, they require fine-tuning for highly specific tasks like sequential recommendation. An LLM should be ideally fully fine-tuned (Kang et al., 2023; Whitehouse et al., 2023) to learn to comprehend domain-specific flattened text sequences and the underlying user interest evolution. Similar to most sequence modeling approaches like SASRec (Kang and McAuley, 2018), BERT4REC (Sun et al., 2019), Recformer (Li et al., 2023c), and GPT4Rec (Li et al., 2023d), we adopt next-item prediction as the sequential recommendation task.
Based on sentence-completion (Li et al., 2023d) or question-answering  (Zhang et al., 2023) styled prompting, our template design is further inspired by the idea of few-shot learning (Brown et al., 2020). Specifically, we prefix user sequence and each item in the user sequence with distinctive text indicators that 1) allow the model to better understand data pattern and text format and 2) increase the LLM response coherence of the input sentence. We further propose a Contrastive Aligned Generative LLM Recommendation (CALRec) framework to adapt an LLM for sequential recommendation, inspired by the effectiveness of contrastive learning (CL) in mapping-based RecSys (Hou et al., 2022; Li et al., 2023c) and CL research in NLP and other areas  (Gao et al., 2021; Li et al., 2023a). CALRec features the following:
@@
Similar to previous work (Kang and McAuley, 2018; Sun et al., 2019; Li et al., 2023c), we leverage the Amazon Review dataset (Ni et al., 2019) and compare our results to open-source state-of-the-art (SotA) baselines. We observed a non-negligible percentage of consecutive identical activities by the same users in the raw datasets which might trivialize the next-item prediction task. We thus train and evaluate the model on deduplicated sequences and propose to adopt ‘Last Item Repeater’, a training-free, text-statistic baseline for model comparison (§4 and §6). We also revisit the existing evaluation metrics of the previous works, and propose to report both optimistic and pessimistic metric scores for text-based RecSys (§4). Through comprehensive experiments, we discover that our CALRec method demonstrates superior performance and outperforms existing SotA by a considerable margin in most evaluation metrics. A systematic ablation study indicates the effectiveness of the two-stage training paradigm, our contrastive alignment, and our template design among the various data categories examined in our experiments.
@@ 2. Related Work
Traditional recommender systems like Collaborative Filtering (Goldberg et al., 1992) learn relevance through user-item co-occurrence. Early collaborative filtering approaches like Matrix Factorization (MF) (Koren et al., 2009) are ID-only where both user and item IDs are embedded in a shared embedding space. Later MF improvements incorporated other attributes (Shi et al., 2014) such as time and location. Nevertheless, these approaches did not capture the sequential nature of user behavior which gave rise to a class of Sequential Recommenders. The initial Sequential Recommenders, similar to MF, were ID-only models, exemplified by GRU4Rec (Hidasi et al., 2016), SASRec (Kang and McAuley, 2018), and Bert4Rec (Sun et al., 2019). Subsequent research expanded these models by integrating additional attributes beyond IDs, leading to the development of methods such as FDSA (Zhang et al., 2019), S3-Rec (Zhou et al., 2020), and UniSRec (Hou et al., 2022). Due to the challenges associated with IDs in real world Recommender Systems (Shi et al., 2014), researchers explored removing them in favor of attribute-only models (Hou et al., 2022; Li et al., 2023c). However, this shift led to quality issues, as the valuable ‘memorization characteristics’ provided by IDs were lost. Another thread of research explored deriving IDs from attribute embedding (Rajput et al., 2023; Singh et al., 2023; Hou et al., 2023), which mitigates some of the challenges associated with random IDs, albeit at the cost of increased system complexity.
@@
With the advent of pretrained LLMs, which include encoder-only models like BERT (Devlin et al., 2019; Li et al., 2023c), encoder-decoder models like T5 (Raffel et al., 2020), and decoder-only models like PaLM (Chowdhery et al., 2023; Anil et al., 2023), GPT (Brown et al., 2020), and LLama (Touvron et al., 2023), researchers explored leveraging these backbones in sequential recommendation tasks, harnessing the extensive ‘knowledge’ encoded within their parameters. Some approaches utilized zero-shot or few-shot methods (Wang and Lim, 2023; Liu et al., 2023; Hou et al., 2024), while others extract item text embeddings from LLMs, both with frozen LLM parameters.
@@
Previous work also explored fine-tuning encoder-only, encoder-decoder or decoder-only LLMs to incorporate recommendation knowledge into the LLMs (Li et al., 2023c; Raffel et al., 2020; Wu et al., 2023; Lin et al., 2023; Kang et al., 2023; Zhang et al., 2023; Bao et al., 2023; Ji et al., 2024; Zhu et al., 2024). P5 (Geng et al., 2022) fine-tuned a pre-trained T5 (Raffel et al., 2020) model to solve various common RecSys tasks, including sequential recommendation. In contrast to P5, which frames the sequential recommendation task as an ID generation problem, our approach focuses on generating item attribute text and features a two-tower training framework, enhancing model’s understanding of items with their attributes within the context of user-item interaction sequences. GPT4Rec (Li et al., 2023d) fine-tunes a pre-trained GPT-2 model to generate the next-item titles. Similar to GPT4Rec, we use BM25 (Robertson et al., 2009) to search for items in the item corpus. However, unlike GPT4Rec which was directly fine-tuned for next-item generation as a sentence completion task on each target domain respectively, we are among the first to explore using the contrastive loss as an auxiliary objective to fine-tune an autoregressive LLM on the sequential recommendation task, and our work also features a multi-stage fine-tuning framework, enabling our model to generalize to multiple target domains. We include the list of baseline models we compare against and their details in §4.
@@ 3.1. Preliminaries and Terminology
We assume a set of users $\mathcal{U}$, where a user $u\in\mathcal{U}$ is associated with an interaction sequence comprising $n$ items $(I_{1},I_{2},...,I_{n})$ arranged in chronological order (note that the value of $n$ may vary across different users). The sequential recommendation task is then defined as follows: given the user’s historical interaction sequence $I_{1:n-1}=(I_{1},I_{2},...,I_{n-1})$, a sequential recommender aims to predict the target item $I_{n}$, the last item in the user’s full interaction sequence. In our study, we use pure text to represent each item and user interaction sequence (§3.2), and we adapt autoregressive LLMs pretrained on large corpora of text for sequential recommendation. Specifically, we pose the task as a sequence-to-sequence generation task where the input is the text description of $I_{1:n-1}$ and the expected output is that of $I_{n}$. Our primary motivation for formulating the problem entirely within the text domain is to exploit the rich textual information inherent in real-world data and to capitalize on LLMs’ strong language understanding and reasoning capabilities.
@@ 3.2. Data Format and Template Design
Different from our method, many previous text-based approaches do not specify an item prefix and simply leverage a one-off user history suffix (Li et al., 2023d; Zhang et al., 2023) (some approaches even do not specify any user suffix at all (Li et al., 2023c)).
@@ 3.3. Two-Stage Fine-Tuning
Previous work (Hou et al., 2022; Li et al., 2023c) indicates that pretraining a RecSys on a large volume of data (various data categories) provides favorable inductive bias for downstream tasks (a specific and smaller-scale data category for evaluation after further fine-tuning on them). Analogously, based on an LLM already pretrained on large, diverse text corpora, we propose a two-stage fine-tuning framework that consists of Multi-Category Joint Fine-Tuning (Stage I) and Category-Specific Fine-Tuning (Stage II): the latter benefits from the former as a result of transfer learning. Our Stage I fine-tuning allows the model to adapt to the sequential recommendation problem setting and learn data patterns in a category-agnostic fashion. Inspired by Xue et al. (2021), we adopt a sampling rate for each data category proportional to $|\mathcal{U}|^{0.3}$, where $|\mathcal{U}|$ is the size (number of users) of each category, for the purpose of combating the data imbalance issue, preventing a few extra-large categories from dominating the Stage I fine-tuning set. After deriving the multi-category fine-tuned model, we further adapt the model to each category respectively via Category-Specific Fine-Tuning (Stage II). In both stages, the training objective is the same, introduced in §3.4.
@@ 3.4. Training Objectives
Next-Item Generation Objective. The primary objective employed for LLM fine-tuning is the Next-Item (Text) Generation (NIG) Objective. NIG aims to generate the text description of the target item given the text description of the previous items in a user’s history. Denote the text-tokenized user sequence (for a user) as $\textsc{t}=(t_{1},t_{2},...,t_{l})$, where $l$ is the tokenized sequence length. The first $m$ tokens belong to all the items in the user sequence except for the last item (i.e., the target item), followed by $l-m$ tokens for the target item. We propose this next-item generation objective to adapt an LLM for sequential recommendation:
(1) $\mathcal{L}_{NIG}=\mathop{\mathbb{E}}\limits_{\textsc{t}\sim p(\textsc{t})}\left[\sum_{j=m+1}^{l}\text{log}\,P(t_{j}|t_{1:j-1};\theta)\right]$
where p(t) denotes the data distribution, and $\theta$ is the whole set of trainable parameters of the chosen LLM.
Auxiliary Contrastive Alignment. One shortcoming of the NIG is that it deals with the input and output on the token level rather than on the item/user level. Therefore, in addition to the token-level NIG objective, we also investigate auxiliary contrastive objectives that operate on the item/user level. Specifically, we adopt a two-tower training framework where one tower only takes in the target item, and the mean-pooled feature in the final hidden layer is denoted as $\mathbf{v}^{T}$. The second tower takes the whole user sequence and similarly we denote the mean-pooled feature corresponding to the user history sequence as $\mathbf{v}^{U}$ and the feature for the target item conditioned on the user history as $\mathbf{v}^{T|U}$ (see also Fig. 1). We experiment with two contrastive losses for user- and item-level alignments following the InfoNCE loss  (Oord et al., 2018), a standard and robust choice in contrastive learning research  (Chen et al., 2020; Gao et al., 2021; Li et al., 2022, 2023a). Below, we present the calculation of these losses on a batch of data:
@@
Our final training objective is a mixture of $\mathcal{L}_{NIG}$ and contrastive losses as shown in Fig. 1: